OpenStudy (anonymous):

Find the unit rate 26 points/ 3 quarters

jimthompson5910 (jim_thompson5910):

Divide each component by 3 to get 26/3 = 8.667 and 3/3 = 1 So 26 points/ 3 quarters = 8.667 points/1 quarter

OpenStudy (anonymous):

The anwser is 8\[the anwser is 8 2/3 points /1quater\]

OpenStudy (anonymous):

the anwser is 8 and 2/3 over 3 quarters

jimthompson5910 (jim_thompson5910):

ah so they want a mixed number 26/3 = (24+2)/3 = 24/3 + 2/3 = 8 + 2/3 = 8 2/3 This means that 26/3 = 8 2/3 So the final answer is then 8 2/3 points over 1 quarter

OpenStudy (anonymous):

Show how I found this

jimthompson5910 (jim_thompson5910):

see above

jimthompson5910 (jim_thompson5910):

oh and you're still dividing the bottom component 3 by 3 to get 3/3 = 1, so that doesn't change

OpenStudy (anonymous):

Thank you so much....I jsut started learning this and we only had like 1/3 a bell to learn it because we had to go to curch in class.

jimthompson5910 (jim_thompson5910):

well you'll get the hang of it with more practice don't worry

OpenStudy (anonymous):

not 1/3 i ment 1/2 a bell

OpenStudy (anonymous):

Thanks for the support :)

jimthompson5910 (jim_thompson5910):

yw